# Towards Optimizing the Expected Performance of Sampling-Based Quantum-Inspired Algorithms

### v2

@@ 5.2 Sample efficiency for general well-conditioned states
Definition 3 (Well-conditioned state [13]).
@@
Stabilizer states (e.g., the GHZ state), the W state, and the Dicke state with $k$ excitations [25, 3] are well-conditioned with $\alpha=1$, $\alpha=1/n$, and $\alpha=O(1/n^{k})$, respectively.
@@ 7 Conclusion
Appendix A Proof of Proposition 3
@@
We use the importance sampling technique [16, 23, 26]. Define a random variable
@@
$\mathbb{E}[Z(i)]_{i\sim\mathcal{D}^{(p)}_{x}}$ $=\sum_{i=1}^{n}\frac{|x_{i}|^{p}}{||x||_{p}^{p}}\text{sgn}(x_{i})|x_{i}|^{1-p}y_{i}$
$=\frac{\sum_{i=1}^{n}x_{i}y_{i}}{||x||_{p}^{p}}$
$=\frac{\langle x,y\rangle}{||x||_{p}^{p}}$
@@
$=\sum_{i=1}^{n}\frac{|x_{i}|^{p}}{||x||_{p}^{p}}|x_{i}|^{2-2p}y_{i}^{2}$
$=\frac{\sum_{i=1}^{n}|x_{i}|^{2-p}y_{i}^{2}}{||x||_{p}^{p}}$
$=\frac{\langle x^{(2-p)},y^{(2)}\rangle}{||x||_{p}^{p}}.$
Therefore, the random variable $||x||_{p}^{p}Z(i)$ has mean $\langle x,y\rangle$ and standard deviation less than or equal to $||x||_{p}^{p/2}\sqrt{\langle x^{(2-p)},y^{(2)}\rangle}$. It is known that the median of $6\log\frac{1}{\delta}$ sample means with each sample containing $\frac{9}{2\epsilon^{2}}$ data points is $\epsilon\sigma$-close to $\mu$ with probability at least $1-\delta$, where $\mu$ and $\sigma$ are the mean and variance of the random variable, respectively. The result directly follows. ∎
Appendix B Proof of Proposition 4
@@
i.e., $g(\hat{p})=g(2-\hat{p})$, which is a contradiction. Hence, $f(p)$ attains its minimum at $p=1$. ∎
Now we prove Proposition 4.
@@
Let $\tilde{\sigma}=||x||_{p}^{p/2}\sqrt{\langle x^{(2-p)},y^{(2)}\rangle}$ and $\mathbb{E}_{y}[y_{i}^{2}]=\upsilon$, where $\upsilon$ is a constant. If we fix $\epsilon\tilde{\sigma}=c_{1}$ and $1-\delta=c_{2}$ in Proposition 3, then the number of queries and samples is $O\left(\frac{\tilde{\sigma}^{2}}{c_{1}^{2}}\log\frac{1}{1-c_{2}}\right)=O\left(\frac{||x||_{p}^{p}\langle x^{(2-p)},y^{(2)}\rangle}{c_{1}^{2}}\log\frac{1}{1-c_{2}}\right)$. Then from Proposition 3, we see that reducing $||x||_{p}^{p}||x||_{2-p}^{2-p}$ with respect to $p$ allows us to achieve the same precision with smaller number of queries and samples on average. This is because
$\mathbb{E}_{y}\big{[}\text{Var}[||x||_{p}^{p}Z(i)]_{i\sim\mathcal{D}^{(p)}_{x}}\big{]}$ $=||x||_{p}^{p}\mathbb{E}_{y}[\langle x^{(2-p)},y^{(2)}\rangle]-\mathbb{E}_{y}[\langle x,y\rangle^{2}]$
$=||x||_{p}^{p}\langle x^{(2-p)},\mathbb{E}_{y}[y^{(2)}]\rangle-\mathbb{E}_{y}[\langle x,y\rangle^{2}]$
$=\upsilon||x||_{p}^{p}||x||_{2-p}^{2-p}-\mathbb{E}_{y}[\langle x,y\rangle^{2}]$
@@
Appendix C Proof of Corollary 2
@@
$\mathbb{E}_{x}\left[\upsilon||x||_{p}^{p}||x||_{2-p}^{2-p}-\mathbb{E}_{y}[\langle x,y\rangle^{2}]\right]=\mathbb{E}_{x}\left[\upsilon\sum_{i=1}^{n}|x_{i}|^{p}\sum_{i=1}^{n}|x_{i}|^{2-p}-\mathbb{E}_{y}\left[\left(\sum_{i=1}^{n}x_{i}y_{i}\right)^{2}\right]\right].$
@@
Appendix D Proof of Proposition 5
@@
$\text{SQ}_{p}(A)$ implies query access to $||A^{(j)}||_{p}^{p}$. In the form of rejection sampling, $P$ is the distribution over $i\in[m]$ formed by first sampling $j\in[n]$ with probability proportional to $||x_{j}A^{(j)}||_{p}^{p}$ and then sampling $i\in[m]$ from $\mathcal{D}_{A^{(j)}}^{(p)}$. The target distribution is $Q:=\mathcal{D}_{Ax}^{(p)}$. Note that
$P(i)$ $=\sum_{j=1}^{n}\frac{||x_{j}A^{(j)}||_{p}^{p}}{\sum_{j^{\prime}=1}^{n}||x_{j^{\prime}}A^{(j^{\prime})}||_{p}^{p}}\frac{|A_{ij}|^{p}}{||A^{(j)}||_{p}^{p}}$
$=\frac{\sum_{j=1}^{n}|x_{j}A_{ij}|^{p}}{\sum_{j=1}^{n}||x_{j}A^{(j)}||_{p}^{p}}.$
@@
$=\frac{\sum_{j=1}^{n}||x_{j}A^{(j)}||_{p}^{p}}{\sum_{j=1}^{n}|x_{j}A_{ij}|^{p}}\frac{|(Ax)_{i}|^{p}}{||Ax||_{p}^{p}}\frac{1}{r_{i}}.$
The terms that do not depend on $i$ are $\sum_{j=1}^{n}||x_{j}A^{(j)}||_{p}^{p}$ and $||Ax||_{p}^{p}$. Therefore, to make $M$ constant,
@@
$n^{p-1}\sum_{j=1}^{n}|x_{j}A_{ij}|^{p}$ =$=\left\{\left(\sum_{j=1}^{n}1^{\frac{p}{p-1}}\right)^{\frac{p-1}{p}}\left(\sum_{j=1}^{n}|x_{j}A_{ij}|^{p}\right)^{\frac{1}{p}}\right\}^{p}$
@@
for $p>1$ (the inequality trivially holds for $p=1$; we just can’t place $p-1$ in the denominator). To make $r_{i}\leq 1$, we choose $k=1/n^{p-1}$ which gives $M=n^{p-1}\sum_{j=1}^{n}||x_{j}A^{(j)}||_{p}^{p}/||Ax||_{p}^{p}$. Note that in each iteration, the dominant time complexity is that for sampling $j\in[n]$ and calculating $r_{i}$, which requires $O(n)$ queries. Since $M$ is the expected number of iterations for a single valid sample, we have our result. ∎
Appendix E Proof of Observation 1
@@
From (D), it immediately follows that $n^{p-1}\sum_{i=1}^{m}\sum_{j=1}^{n}|x_{j}A_{ij}|^{p}\geq\sum_{i=1}^{m}|\sum_{j=1}^{n}x_{j}A_{ij}|^{p}$ and $M(p)\geq 1$. If $\text{sgn}(A_{i})$=±$=\pm\text{sgn}(x)$, then $|\sum_{j=1}^{n}x_{j}A_{ij}|=\big{|}\sum_{j=1}^{n}\pm|x_{j}||A_{ij}|\big{|}=\sum_{j=1}^{n}|x_{j}||A_{ij}|$. Therefore, $M(1)=1\leq$$M(p)$. Now consider the extreme case: $A^{(1)}$ is nonzero, $A^{(j)}$ is a zero vector for each $j\in[n]\setminus\{1\}$, and $x_{1}$ is nonzero. We have $M(p)=n^{p-1}$ and if $p>1$, then $M(p)/M(1)=n^{p-1}$ can get arbitrarily large with increasing $n$. ∎
Appendix F Proof of Lemma 2
@@
$\text{Cov}\left(\sum_{j=1}^{n}x_{j}A_{ij},\sum_{j=1}^{n}x_{j}A_{i^{\prime}j}\right)$ $=\mathbb{E}\left[\sum_{j,j^{\prime}=1}^{n}x_{j}x_{j^{\prime}}A_{ij}A_{i^{\prime}j^{\prime}}\right]-\mathbb{E}\left[\sum_{j=1}^{n}x_{j}A_{ij}\right]\mathbb{E}\left[\sum_{j=1}^{n}x_{j}A_{i^{\prime}j}\right]$
@@
converges in distribution to $\mathcal{N}(0,(\alpha^{2}+\beta^{2})\sigma_{f}^{4})$ as $n\rightarrow\infty$. This implies that the joint distribution of $\sum_{j=1}^{n}x_{j}A_{ij}/\sqrt{n}$ and $\sum_{j=1}^{n}x_{j}A_{i^{\prime}j}/\sqrt{n}$ converges to the bivariate normal distribution $\mathcal{N}\left(\begin{pmatrix}0\\ 0\\ \end{pmatrix},\begin{pmatrix}\sigma_{f}^{4}&0\\ 0&\sigma_{f}^{4}\\ \end{pmatrix}\right)$, i.e., the two random variables become independent Gaussians. Consequently, the absolute moments $|\sum_{j=1}^{n}x_{j}A_{ij}/\sqrt{n}|^{p}$ and $|\sum_{j=1}^{n}x_{j}A_{i^{\prime}j}/\sqrt{n}|^{p}$ also become independent. Averaging over $i$, we conclude that as $m,n\rightarrow\infty$, the variance of $\sum_{i=1}^{m}|\sum_{j=1}^{n}x_{j}A_{ij}/\sqrt{n}|^{p}/m$ converges to 00 and $\sum_{i=1}^{m}|\sum_{j=1}^{n}x_{j}A_{ij}/\sqrt{n}|^{p}/m$ converges in probability to $\tilde{\mu}_{f,p}$ due to Theorem 9. ∎
Appendix G Plots for Lemma 2
@@ Appendix I Derivation of (9)
From Hoeffding’s inequality, it can be shown that $F(\rho,\sigma)$ lies in the range $[\tilde{Y}-2\epsilon,\tilde{Y}+2\epsilon]$ with probability $\geq 1-2\delta$, where the upper bound on the total number of measurements is
@@
From Hoeffding’s inequality, it can be shown that $F(\rho,\sigma)$ lies in the range $[\tilde{Y}-2\epsilon,\tilde{Y}+2\epsilon]$ with probability $\geq 1-2\delta$, where the upper bound on the total number of measurements is

### v3

@@ 5.2 Sample efficiency for general well-conditioned states
Definition 3 (Well-conditioned state [flammia2011direct]).
@@
Stabilizer states (e.g., the GHZ state), the W state, and the Dicke state with $k$ excitations [kiesel2007experimental, bartschi2019deterministic] are well-conditioned with $\alpha=1$, $\alpha=1/n$, and $\alpha=O(1/n^{k})$, respectively.
@@ 7 Conclusion
Appendix A Proof of Observation 1
@@
We use the importance sampling technique [goertzel1949quota, kahn1951estimation, kloek1978bayesian]. Define a random variable
@@
[…]
@@
[…]
Therefore, the random variable $x||_{p}^{p}Z(i)$ has mean $\langle x,y\rangle$ and standard deviation less than or equal to $x||_{p}^{p/2}\sqrt{\langle x^{(2-p)},y^{(2)}\rangle}$. It is known that the median of $6\log\frac{1}{\delta}$ sample means with each sample containing $\frac{9}{2\epsilon^{2}}$ data points is $\epsilon\sigma$-close to $\mu$ with probability at least $1-\delta$, where $\mu$ and $\sigma$ are the mean and variance of the random variable, respectively. The result directly follows. ∎
Appendix B Proof of Observation 2
@@
i.e., g(p^)=g(2p^)g($\hat${p})=g(2-$\hat${p}), which is a contradiction. Hence, $f(p)$ attains its minimum at $p=1$. ∎
Now we prove Observation 2.
@@
[…]
@@
Appendix C Proof of Observation 3
@@
[…]
@@
Appendix D Proof of Observation 4
@@
$\text{SQ}_{p}(A)$ implies query access to $A^{(j)}||_{p}^{p}$. In the form of rejection sampling, $P$ is the distribution over $i\in[m]$ formed by first sampling $j\in[n]$ with probability proportional to $x_{j}A^{(j)}||_{p}^{p}$ and then sampling $i\in[m]$ from $\mathcal{D}_{A^{(j)}}^{(p)}$. The target distribution is $Q:=\mathcal{D}_{Ax}^{(p)}$. Note that
[…]
@@
[…]
@@
$n^{p-1}\sum_{j=1}^{n}|x_{j}A_{ij}|^{p}$ $=\left\{\left(\sum_{j=1}^{n}1^{\frac{p}{p-1}}\right)^{\frac{p-1}{p}}\left(\sum_{j=1}^{n}|x_{j}A_{ij}|^{p}\right)^{\frac{1}{p}}\right\}^{p}$
@@
[…]
Appendix E Proof of Observation 6
@@
[…]
Appendix F Proof of Observation 7
@@
$\text{Cov}\left(\sum_{j=1}^{n}x_{j}A_{ij},\sum_{j=1}^{n}x_{j}A_{i^{\prime}j}\right)$
$=$ $\mathbb{E}\left[\sum_{j,j^{\prime}=1}^{n}x_{j}x_{j^{\prime}}A_{ij}A_{i^{\prime}j^{\prime}}\right]-\mathbb{E}\left[\sum_{j=1}^{n}x_{j}A_{ij}\right]\mathbb{E}\left[\sum_{j=1}^{n}x_{j}A_{i^{\prime}j}\right]$
@@
converges in distribution to $\mathcal{N}(0,(\alpha^{2}+\beta^{2})\sigma_{f}^{4})$ as $n\rightarrow\infty$. This implies that the joint distribution of $\sum_{j=1}^{n}x_{j}A_{ij}/\sqrt{n}$ and $\sum_{j=1}^{n}x_{j}A_{i^{\prime}j}/\sqrt{n}$ converges to the bivariate normal distribution $\mathcal{N}\left(\begin{pmatrix}0\\ 0\\ \end{pmatrix},\begin{pmatrix}\sigma_{f}^{4}&0\\ 0&\sigma_{f}^{4}\\ \end{pmatrix}\right)$, i.e., the two random variables become independent Gaussians. Consequently, the absolute moments $|\sum_{j=1}^{n}x_{j}A_{ij}/\sqrt{n}|^{p}$ and $|\sum_{j=1}^{n}x_{j}A_{i^{\prime}j}/\sqrt{n}|^{p}$ also become independent. Averaging over $i$, we conclude that as $m,n\rightarrow\infty$, the variance of $\sum_{i=1}^{m}|\sum_{j=1}^{n}x_{j}A_{ij}/\sqrt{n}|^{p}/m$ converges to 0 and $\sum_{i=1}^{m}|\sum_{j=1}^{n}x_{j}A_{ij}/\sqrt{n}|^{p}/m$ converges in probability to $\tilde{\mu}_{f,p}$ due to Theorem 5. ∎
Appendix G Plots for Observation 7
@@ Appendix I Derivation of (9)
From Hoeffding’s inequality, it can be shown that $F(\rho,\sigma)$ lies in the range $\tilde{Y}-2\epsilon,\tilde{Y}+2\epsilon]$ with probability $\geq 1-2\delta$, where the upper bound on the total number of measurements is
@@
From Hoeffding’s inequality, it can be shown that $F(\rho,\sigma)$ lies in the range $\tilde{Y}-2\epsilon,\tilde{Y}+2\epsilon]$ with probability $\geq 1-2\delta$, where the upper bound on the total number of measurements is